# Part I: Tricks or Traps? A Deep Dive into RL for LLM Reasoning

### v1

@@ 4.1.2 Impact of the standard deviation term in advantage normalization
To determine whether the calculation method of the standard deviation is the key module causing the difference in normalization performance, we employ the batch-level calculation, which exhibited unstable performance in the previous section, to calculate the mean of advantage, and conduct ablation experiments on the standard deviation term. This can be formalized as:
@@
We separately recorded the accuracy after training on simple and difficult data. The curves of easy data in Figure 6 show that the policy rapidly converges to highly consistent behaviors, leading to a highly concentrated distribution of reward values. Correspondingly, the standard deviation of the reward distribution swiftly declines to a low value. Applying standard deviation-based normalization in this setting results in an exceedingly small denominator, which excessively amplifies reward and advantage values. This, in turn, induces abnormally large gradients, destabilizes training, and can even trigger gradient explosions. Therefore, these experimental results empirically verify our conjecture that the standard deviation term is the key mechanism for the advantage normalization.
@@ 4.1.3 Reconstruct a robust normalization technique
Section 4.1.2 highlights the critical role of the standard deviation in determining the effectiveness of the advantage normalization mechanism. This leads to the final requirement: Is there a more robust and effective combination of mean and standard deviation for reward shaping? To explore this, we adopted the stable group-level mean calculation method demonstrated in section 4.1.1, paired with two approaches for computing the standard deviation: local (group-level) and global (batch-level). We then evaluated the performance of these combinations across two model sizes.
@@ 4.2 Clip-Higher
$\varepsilon_{high}$ denotes the upper bound of the Clip mechanism and $\varepsilon_{low}$ represents the lower bound. Unlike the original clip that enforces proportional fairness, Clip-Higher introduces a higher upper bound for advantage, giving low-probability tokens more improving space. By expanding exploration potential in low-probability regions, this technique effectively mitigates entropy collapse. However, the lack of in-depth analysis of the underlying mechanism and the absence of detailed usage guidelines have left practitioners confused about the appropriate scenarios for using Clip-Higher, as well as the ideal upper bound settings under different conditions. In this section, we address the aforementioned remaining issues through a series of comprehensive experiments.
@@ 4.2.1 In which settings should we clip higher
Through extensive empirical practice, we observe that the advantage clip technique demonstrates distinct effectiveness across different model architectures. To examine this, this section employs the non-aligned (base) model and the aligned (instruct) model with various sizes to clearly demonstrate the sensitivity of the Clip mechanism, summarize the usage guidelines for Clip higher from a modeling perspective.
@@ 4.2.2 Analyzing the effectiveness of Clip-Higher from a linguistic perspective
Building on our token-level demonstration of Clip-Higher’s behavior in section 4.2.1, we now analyze its impact on reasoning logic through token-level linguistics. As illustrated in Figure 11, setting an upper bound to $0.2$ imposes stringent constraints on policy updates by limiting substantial probability deviations for individual tokens. Under these stricter conditions, our analysis reveals that clipping predominantly affects connective tokens such as “therefore”, “if”, and “but”. These tokens frequently appear at the beginnings of sentences, serving as key semantic markers or transition words within dialog generation. Such connectors often introduce new directions in reasoning. However, their probability ratios between updated and old policies frequently exceed clipping thresholds, triggering aggressive suppression in PPO optimization. While this traditional clipping ensures stability in the overall token distribution, it may restrict the model’s capacity to generate innovative or diverse argumentative reasoning structures by constraining flexibility in the use of discourse-level connectives.
@@ 4.3 Loss Aggregation
The strategy of loss aggregation directly determines the contribution of each sample or token to the overall gradient during optimization (Liu et al., 2025b). Common strategies include token-level and sequence-level aggregation. The sequence-level aggregation adopted by GRPO (Shao et al., 2024) first averages the loss across all tokens within each sample, then averages these per-response losses across the batch, thereby assigning equal weight to each response regardless of its length. However, Yu et al. (2025) highlights a flaw in this method: longer responses possess a diminished influence per token on the total loss, hindering the model’s ability to learn effectively from diverse quality reasoning in lengthier responses. This can reduce the model’s capacity to learn from long, complex answers, and may bias optimization toward brevity, since shorter correct responses receive larger gradient updates, while longer incorrect responses are insufficiently penalized (Liu et al., 2025a).
@@
In response to this issue, Yu et al. (2025) turns to a token-level calculation approach. Here, losses are calculated by summing the loss across all tokens from all samples and then normalizing by the total token count, guaranteeing an equal contribution from each token regardless of response length. Despite the widespread adoption of these methods, existing analyses remain trivial. In this section, we provide a detailed empirical comparison of the two loss calculation techniques across diverse training data distributions. The evaluation comprehensively assesses the effectiveness of these methods from the perspective of model type.
@@ 4.3.1 Does token-level loss aggregation suit all settings?
To systematically evaluate the effectiveness of different loss aggregation strategies, we compare token-level and sequence-level loss aggregation on both base and aligned versions of Qwen3-8B, as shown in Figures 13 and 18. For base models, token-level loss consistently improves convergence, peak accuracy, and robustness by ensuring each token contributes equally to the optimization signal, especially on challenging datasets. However, as illustrated in Figure 13 (bottom 2 rows), this advantage does not show in aligned models. In fact, sequence-level aggregation outperforms token-level loss across most datasets and settings, both in convergence speed and final accuracy. Further analysis reveals that aligned models already possess strong and stable reasoning, making the equalization of token-level gradients unnecessary or even detrimental. In these cases, sequence-level aggregation better preserves the structure and consistency of high-quality, aligned outputs.
@@ 4.4 Overlong Filtering
During the training of LLMs, a fixed maximum generation length is often set for truncation to ensure training efficiency and save computational costs (Chen et al., 2025; Team et al., 2025). However, recent studies have revealed that in more complex reasoning tasks, this strategy can prematurely end multi-step tail reasoning processes, particularly noticeable in the early training stages. Consequently, coherent and well-structured reasoning is often cut short before reaching the final answer, causing them to be falsely labeled as negative samples by the model. This noise, akin to penalties, can contaminate the training signal, reducing sample utilization efficiency and learning effectiveness.
@@ 4.4.1 When to use the overlong filtering
To further investigate this effect, Figure 15 (Left) shows the distribution of filtered responses exceeding the maximum length. Notably, in the 20k setting, both correct and incorrect samples are filtered more frequently due to repetitive or non-terminating outputs, a hallmark of degenerate generation. This indicates that, with higher length limits, the overlong filtering strategy primarily filters out unproductive or “negative” samples that contribute little to model learning. As illustrated in Figure 15 (Right), we observed that during RL training on models fine-tuned with instructions, the proportion of “repetitive but unable to terminate normally” samples within the overall set of overlong samples gradually increased as training progressed. This indicates a degradation in the model’s ability to accurately model end-of-sequence (EOS) tokens, leading to behavioral defects in the inference stage, such as output redundancy and hard in terminating the generation. After introducing the overlong filtering mechanism, the proportion of abnormal samples that are “repetitive but unable to terminate” significantly decreased during training. This shift suggests that the model can more accurately distinguish between “completed generation” and “truncated generation” samples during training, effectively avoiding invalid learning from truncated samples. Furthermore, this mechanism may unlock the policies’ ability to accurately model termination behaviors during generation, enabling them to appropriately ignore unfinished inference samples, rather than mistakenly penalizing them as negative examples.
@@
Figure 16: Test accuracy of non-aligned models trained via three RL methods, i.e., Lite PPO (ours), GRPO (Shao et al., 2024) and DAPO (Yu et al., 2025).
@@ 5 A simple combination: Lite PPO
Building on the in-depth mechanism analysis and empirical evaluations presented in previous sections, we derive two key technique guidelines for non-aligned models: (i) For small and medium-sized non-aligned models, i.e., 4B-Base and 8B-Base, the technique that can provide significant performance improvement is the advantage normalization introduced in section 4.1.3. This technique shapes sparse rewards into more robust guiding signals through group-level mean calculation and batch-level standard deviation calculation. (ii) Token-level loss aggregation emerges as another highly effective technique for non-aligned models, with Section 4.3.1 experiments demonstrating its particular efficacy for base model architectures.
We therefore propose the following empirically motivated hypothesis: Given the individually superior performance of advantage normalization (group-level mean, batch-level std) and token-level loss aggregation over alternative techniques, their synergistic combination should show robust improvements in policy optimization. To validate this, we integrate both techniques, called Lite PPO, into non-aligned models that use the vanilla PPO loss without the critic. The results shown in Figure 16 indicate that Lite PPO outperforms the technique-heavy algorithm DAPO, which involves Group-level Normalization, Clip-Higher, Overlong Reward Shaping, Token-level Loss, Dynamic Sampling, and the strong and widely-used RL4LLM algorithm GRPO.
Specifically, in the first two rows of Figure 16, Lite PPO exhibits a stable upward trend on small models lacking basic reasoning ability. In contrast, other policies collapse rapidly after reaching their peak. This significant advantage results from the normalization technique introduced in Takeaway 3, which effectively counters the interference induced by homogeneous reward distributions characteristic of datasets with non-uniform reward levels (easy and hard). We further evaluate Lite PPO on larger base models. As shown in Figure 16, when training 8B-Base models with inherent long-tail generation capabilities on the hard dataset, Lite PPO also demonstrates superior performance. This improvement stems from Lite PPO eliminating overlong filtering (which typically restricts small models’ ability to generate complex long-tail outputs; Takeaway 8), and shifting to token-level loss aggregation (which shows better efficiency on base models; Takeaway 7).
@@ 6 Conclusion
The rapid advancement of reinforcement learning (RL) in enhancing large language models (LLMs) has ushered in a transformative era for complex reasoning tasks. However, the proliferation of RL4LLM research has also introduced significant challenges, including conflicting methodologies and a lack of cohesive guidelines for technique selection. This work addresses these critical issues by conducting a systematic, reproducible evaluation of prominent RL techniques under a unified framework, revealing key insights that resolve existing ambiguities and streamline practical implementation.
By disentangling the theoretical and practical mechanisms of techniques like normalization, clipping, and filtering, our study provides actionable guidelines to demystify their applicability across diverse scenarios. Crucially, we show that simplicity can outperform complexity: a minimalist approach (i.e., Lite PPO) combining only two core techniques achieves superior performance over algorithms cluttered with redundant components. This finding challenges the prevailing trend of over-engineering RL pipelines and underscores the importance of contextual adaptability in technique selection. Our work not only resolves the current fragmentation in RL4LLM practice but also lays a foundation for developing standardized frameworks that balance theoretical rigor with engineering efficiency.
Finally, to ensure experimental fairness, this paper consistently uses the Qwen3 series model for policy initialization. However, conclusions may vary across LLM families due to inherent differences in pre-training processes and architectures. The prevailing trend of model closed-sourcing, often driven by commercial or strategic considerations, significantly impedes model-family-level technical analysis. Therefore, we advocate for increased disclosure of implementation details in future technical reports within the industry. This transparency is crucial to bridge the understanding gap between academia and industry, enabling the community to pool collective insights in artificial intelligence.
@@ 7 Future work
We envision this work as the starting point of a sustained effort to guide the evolution of reinforcement learning for LLMs along principled and empirically grounded trajectories. Our future research will focus on: (1) continue to monitoring and critically evaluating developments in RL4LLM, distilling emerging practices into coherent, evidence-based guidelines for both academic and industrial practitioners; (2) leveraging the proposed ROLL framework to consolidate diverse RL algorithms and optimization strategies into a unified, modular suite, enabling flexible composition and benchmarking within a consistent training infrastructure; (3) continuing to explore streamlined RL algorithms that deliver strong empirical performance with minimal engineering overhead. These directions align with our long-term vision to provide the community with clear and reliable guidance, driving the field toward robust, adaptable, and broadly beneficial progress, while advancing RL4LLM through both algorithmic innovations and comprehensive framework support.
@@ Appendix C Detailed Experimental Results
To further solidify the results in Figure 5, we show in Figure 19 the accuracy achieved using the Qwen3-8B-Base model as the initial model, evaluated across different reward scales with batch-level normalization applied.

### v2

@@ 4.1.1 Impact of the standard deviation term in advantage normalization
To test whether the standard deviation term is the critical factor driving differences in normalization performance, we employ the batch-level calculation, which exhibited unstable performance in the previous section, to calculate the mean of advantage, and conduct ablation experiments on the standard deviation term. This can be formalized as:
@@
We separately recorded the accuracy after training on simple and difficult data. The curves of easy data in Figure 5 show that the policy rapidly converges to highly consistent behaviors, leading to a highly concentrated distribution of reward values. Correspondingly, the standard deviation of the reward distribution swiftly declines to a low value. Applying standard deviation-based normalization in this setting results in an exceedingly small denominator, which excessively amplifies reward and advantage values. This, in turn, induces excessively large gradients, destabilizes training, and may even trigger gradient explosions. Therefore, these experimental results empirically verify our conjecture that the standard deviation term is the key mechanism for the advantage normalization.
@@ 4.1.2 Reconstruct a robust normalization technique
Section 4.1.1 highlights the critical role of the standard deviation in determining the effectiveness of the advantage normalization mechanism. This raises the question: is there a more robust and effective combination of mean and standard deviation for reward shaping? To explore this, we adopt the group-level mean calculation method, paired with two approaches for computing the standard deviation: local (group-level) and global (batch-level). We then evaluated the performance of these combinations across two model sizes.
@@ 4.2 Clip-Higher
$\varepsilon_{high}$ denotes the upper bound of the Clip mechanism and $\varepsilon_{low}$ represents the lower bound. Unlike the original clip that enforces proportional fairness, Clip-Higher introduces a higher upper bound for advantage, allowing low-probability tokens greater opportunity to increase in probability. By expanding exploration potential in low-probability regions, this technique effectively mitigates entropy collapse. However, the lack of in-depth analysis of the underlying mechanism and the absence of detailed usage guidelines have left practitioners confused about the appropriate scenarios for using Clip-Higher, as well as the ideal upper bound settings under different conditions. In this section, we address the aforementioned remaining issues through a series of comprehensive experiments.
@@ 4.2.1 In which settings should we clip higher
Through extensive empirical practice, we observe that the advantage clip technique demonstrates distinct effectiveness across different model architectures. To examine this, this section employs the non-aligned (base) model and the aligned (instruct) model with various sizes to clearly demonstrate the sensitivity of the Clip mechanism, summarize practical guidelines for Clip-Higher from a modeling perspective.
@@ 4.2.2 Analyzing the effectiveness of Clip-Higher from a linguistic perspective
Building on our token-level demonstration of Clip-Higher’s behavior in section 4.2.1, we now analyze its impact on reasoning logic through token-level linguistics. As illustrated in Figure 10, setting an upper bound to $0.2$ imposes stringent constraints on policy updates by limiting substantial probability deviations for individual tokens. Under these stricter conditions, our analysis reveals that clipping predominantly affects connective tokens such as “therefore”, “if”, and “but”. These tokens frequently appear at the beginnings of sentences, serving as key semantic markers or transition words within dialog generation. Such connectors often introduce new directions in reasoning. However, their probability ratios between updated and old policies frequently exceed clipping thresholds, triggering aggressive suppression in PPO optimization. While this traditional clipping ensures stability in the overall token distribution, it may restrict the model’s capacity to generate innovative or diverse argumentative reasoning structures by limiting flexibility in the use of discourse-level connectives.
@@ 4.3 Loss Aggregation
The strategy of loss aggregation directly determines the contribution of each sample or token to the overall gradient during optimization (Liu et al., 2025b). Common strategies include token-level and sequence-level aggregation. The sequence-level aggregation adopted by GRPO (Shao et al., 2024) first averages the loss across all tokens within each sample, then averages these per-response losses across the batch, thereby assigning equal weight to each response regardless of its length. However, Yu et al. (2025) highlights a flaw in this method: longer responses possess a diminished influence per token on the total loss, hindering the model’s ability to learn effectively from longer, complex responses. This can reduce the model’s capacity to learn from long, complex answers, and may bias optimization toward brevity, since shorter correct responses receive larger gradient updates, while longer incorrect responses are insufficiently penalized (Liu et al., 2025a).
@@
In response to this issue, Yu et al. (2025) turns to a token-level calculation approach. Here, losses are calculated by summing the loss across all tokens from all samples and then normalizing by the total token count, guaranteeing an equal contribution from each token regardless of response length. Despite the widespread adoption of these methods, existing analyses remain limited. In this section, we provide a detailed empirical comparison of the two loss calculation techniques across diverse training data distributions. The evaluation comprehensively assesses the effectiveness of these methods from the perspective of model type.
@@ 4.3.1 Does token-level loss aggregation suit all settings?
To systematically evaluate the effectiveness of different loss aggregation strategies, we compare token-level and sequence-level loss aggregation on both base and aligned versions of Qwen3-8B, as shown in Figures 12 and 17. For base models, token-level loss consistently improves convergence, peak accuracy, and robustness by ensuring each token contributes equally to the optimization signal, especially on challenging datasets. However, as illustrated in Figure 12 (bottom 2 rows), this advantage does not show in aligned models. In fact, sequence-level aggregation outperforms token-level loss across most datasets and settings, both in convergence speed and final accuracy. Further analysis reveals that aligned models already possess strong and stable reasoning, making the equalization of token-level gradients unnecessary or even detrimental. In these cases, sequence-level aggregation better preserves the structure and consistency of high-quality, aligned outputs.
@@ 4.4 Overlong Filtering
During the training of LLMs, a fixed maximum generation length is often set for truncation to ensure training efficiency and save computational costs (Chen et al., 2025; Team et al., 2025). However, recent studies have revealed that in more complex reasoning tasks, this strategy can prematurely end multi-step tail reasoning processes, particularly in early training stages. Consequently, coherent and well-structured reasoning is often cut short before reaching the final answer, causing them to be falsely labeled as negative samples by the model. This noise, akin to penalties, can contaminate the training signal, reducing sample utilization efficiency and learning effectiveness.
@@ 4.4.1 When to use the overlong filtering
To further investigate this effect, Figure LABEL:fig:mask_type shows the distribution of clipped responses exceeding the maximum length. Notably, in the 20k setting, both positive and negative samples are clipped more frequently due to repetitive or non-terminating outputs—a hallmark of degenerate generation. This indicates that, with higher length limits, the overlong mask primarily filters out unproductive or "negative" samples that contribute little to model learning.
Conversely, with a stricter 8k mask threshold, the data mask filters out more samples that are long due to extended—but not necessarily degenerate—reasoning. In this setting, the model is incentivized to produce shorter, more concise responses, discouraging excessive verbosity. Thus, in tasks that do not require long-tail reasoning, data mask primarily mitigates performance degradation from unnecessary length. Meanwhile, if practitioners expect that LLMs generate extremely long reasoning paths, data mask may remove pathological outputs without significantly affecting valid reasoning sequences.
As illustrated in Figure LABEL:repeat_ratio, We observed that during RL training on models fine-tuned with instructions, the proportion of "repetitive but unable to terminate normally" samples within the overall set of overlong samples gradually increased as training progressed. This indicates a degradation in the model’s ability to accurately model end-of-sequence (EOS) tokens, thus leading to behavioral defects in the inference stage, such as output redundancy and difficulties in terminating generation. After introducing the overlong mask mechanism, the proportion of abnormal samples that are "repetitive but unable to terminate" significantly decreased and remained stable at a lower level throughout training. This shift indicates that the model can more accurately distinguish between "generation completed" and "truncated" samples during training, effectively avoiding invalid learning on truncated portions. More importantly, this mechanism may unlock the policies’ ability to accurately model termination behaviors during generation, enabling them to appropriately ignore unfinished inference samples, rather than mistakenly penalizing them as negative examples.
@@
Figure 15: Test accuracy of non-aligned models trained with three RL methods, i.e., Lite PPO (ours), GRPO (Shao et al., 2024) and DAPO (Yu et al., 2025).
@@ 5 A simple combination: Lite PPO
Building on the in-depth mechanism analysis and empirical evaluations presented in previous sections, we derive two key technique guidelines for non-aligned models: (i) For small and medium-sized non-aligned models, i.e., 4B-Base and 8B-Base, the most effective technique is the advantage normalization introduced in section 4.1.2. This technique shapes sparse rewards into more robust guiding signals through group-level mean calculation and batch-level standard deviation calculation. (ii) Token-level loss aggregation emerges as another highly effective technique for non-aligned models, with Section 4.3.1 experiments demonstrating its particular efficacy for base model architectures.
We therefore propose the following empirically motivated hypothesis: Given the individually superior performance of advantage normalization (group-level mean, batch-level std) and token-level loss aggregation over alternative techniques, their combination should yield robust improvements in policy optimization. To validate this, we integrate both techniques, called Lite PPO, into non-aligned models that use the vanilla PPO loss without the critic. The results shown in Figure 15 indicate that Lite PPO outperforms the technique-heavy algorithm DAPO, which involves Group-level Normalization, Clip-Higher, Overlong Reward Shaping, Token-level Loss, Dynamic Sampling, and the strong and widely-used RL4LLM algorithm GRPO.
Specifically, in the first two rows of Figure 15, Lite PPO exhibits a stable upward trend on small models lacking basic reasoning ability. In contrast, other policies collapse rapidly after reaching their peak. This significant advantage results from the normalization technique introduced in Takeaway 3, which effectively counters the interference induced by homogeneous reward distributions characteristic of datasets with non-uniform reward levels (easy and hard). We further evaluate Lite PPO on larger base models. As shown in Figure 15, when training 8B-Base models with inherent long-tail generation capabilities on the hard dataset, Lite PPO also demonstrates superior performance. This improvement results from Lite PPO eliminating overlong filtering (which typically restricts small models’ ability to generate complex long-tail outputs; Takeaway 8), and shifting to token-level loss aggregation (which shows better efficiency on base models; Takeaway 7).
@@ 6 Conclusion
The rapid advancement of reinforcement learning (RL) in enhancing large language models (LLMs) has ushered in a transformative era for complex reasoning tasks. However, the proliferation of RL4LLM research has also introduced significant challenges, including conflicting methodologies and a lack of cohesive guidelines for technique selection. This work addresses these issues by conducting a systematic, reproducible evaluation of prominent RL techniques under a unified framework, revealing key insights that resolve existing ambiguities and streamline practical implementation.
By disentangling the theoretical and practical mechanisms of techniques like normalization, clipping, and filtering, our study provides actionable guidelines that clarify their applicability across diverse scenarios. Crucially, we show that simplicity can outperform complexity: a minimalist approach (i.e., Lite PPO) combining only two core techniques, achieves superior performance over algorithms cluttered with redundant components. This finding challenges the prevailing trend of over-engineering RL pipelines and underscores the importance of contextual adaptability in technique selection. Our work not only resolves the current fragmentation in RL4LLM practice but also lays a foundation for developing standardized frameworks that balance theoretical rigor with engineering efficiency.
Finally, to ensure experimental fairness, this paper consistently uses the Qwen3 series model for policy initialization. However, conclusions may vary across LLM families due to inherent differences in pre-training processes and architectures. The prevailing trend of model closed-sourcing, often driven by commercial or strategic considerations, significantly impedes model-family-level technical analysis. Therefore, we advocate for increased disclosure of implementation details in future technical reports within the industry. Such transparency is crucial to bridge the understanding gap between academia and industry, enabling the community to pool collective insights in artificial intelligence.
@@ 7 Future work
We envision this work as the starting point of a sustained effort to guide the evolution of reinforcement learning for LLMs along principled and empirically grounded trajectories. Our future research will focus on: (1) continuing to monitor and critically evaluate developments in RL4LLM, distilling emerging practices into coherent, evidence-based guidelines for both academic and industrial practitioners; (2) leveraging the proposed ROLL framework to consolidate diverse RL algorithms and optimization strategies into a unified, modular suite, enabling flexible composition and benchmarking within a consistent training infrastructure; (3) continuing to explore streamlined RL algorithms that deliver strong empirical performance with minimal engineering overhead. These directions align with our long-term vision to provide the community with clear and reliable guidance, driving the field toward robust, adaptable, and broadly beneficial progress while advancing RL4LLM through both algorithmic innovations and comprehensive framework support.
@@ Appendix C Detailed Experimental Results
To further solidify the results in Figure LABEL:fig:reward_shape_norm_val_4b_easy, we show in Figure 18 the accuracy achieved using the Qwen3-8B-Base model as the initial model, evaluated across different reward scales with batch-level normalization applied.